OpenStudy (anonymous):

The length of a rectangle is 4/5 feet. The width is 5/8 feet. How much greater is the length of the rectangle than the width? a.1/5 ft b.7/40 ft c.3/20 ft d.1/3 ft

OpenStudy (dwelsher):

21.9%

OpenStudy (whpalmer4):

You need to compute length - width, or \[\frac{4}{5}-\frac{5}{8}\]Start by putting the fractions over a common denominator. You can multiply the first fraction by \(\frac{8}8\) and the second fraction by \(\frac{5}5\). Both fractions are equal to 1, so don't change the value of the fractions you multiply them with, but will give you a common denominator after you have done so.

OpenStudy (whpalmer4):

@dwelsher sorry, that isn't correct... if you wanted to find the percentage that the length is greater than the width, the correct computation is \[\frac{\text{length}-\text{width}}{\text{width}}*100\%\]

OpenStudy (whpalmer4):

you must have divided by length instead of width.

OpenStudy (dwelsher):

lets see if I understand. we want to find how much greater the length is than the width. then first we have to subtract the width from the length to calculate the difference of .8-.625=.175. Is this right? How much greater is .175 of .625?=~27% not21.9 must be a simple question to determine the difference? which would be b?7/40=.175 As you can see I didn't divide by anything. If not please get back to me!

OpenStudy (whpalmer4):

.175/.625 = 0.28 In determining "how much greater is .175 of .625" you're doing that division, whether you realize it or not.

OpenStudy (dwelsher):

Exactly.you are correct.thanks for the help!

OpenStudy (okagurao):

CAN I HAVE A DIRECT ANSWER!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! @dwelsher @whpalmer4

OpenStudy (anonymous):

Help! I don't understand this question!

OpenStudy (anonymous):

the answer is 7/40